# DUMP: Automated Distribution-Level Curriculum Learning for RL-based LLM Post-training

### v1

@@ Abstract
11footnotetext: Corresponding author. Email: kuwan@adobe.com, wezhao@adobe.com.
@@ 1 Introduction
Reinforcement learning (RL)-based post-training has emerged as a powerful approach for enhancing the capabilities of large language models (LLMs), particularly in areas requiring structured reasoning, multi-step inference, and task-specific generalization [1, 2, 3, 4, 5]. By leveraging reward signals derived from task performance, human feedback, or domain-specific metrics, RL provides a flexible alternative to supervised fine-tuning. Unlike imitation-based methods that merely mimic reference outputs, RL-based approaches allow models to optimize directly toward behavioral objectives, making them especially effective for boosting model performance on complex reasoning and agentic tasks.
While RL-based post-training has become a key technique for enhancing LLM capabilities in reasoning, alignment, and coding, one foundational challenge remains underexplored: how to dynamically schedule training across heterogeneous data distributions. In practice, LLMs are post-trained on datasets drawn from a wide variety of sources—ranging from factual QA to math problems and coding tasks—each differing in knowledge/capability relevance, and learning difficulty [6, 7, 8]. This heterogeneity is evident in large-scale post-training datasets such as Tülu 3 [8], where prompts span general dialogue, logic puzzles, STEM problems, and multilingual instructions, with widely varying counts, formats, and alignment objectives. More recently, next-generation post-training pipelines (e.g., Seed-Thinking v1.5 [9]) have shifted toward synthetic data generation with controllable parameters—e.g., configuring logical puzzle difficulty. This allows fine-grained control over the data distribution, making distribution-level curriculum learning both feasible and increasingly important. Despite this, most RL-based pipelines still treat all data distributions equally—uniformly sampling tasks throughout training or relying on static, hand-designed curricula. This static treatment ignores the model’s evolving learning needs and underutilizes the training budget. As reinforcement learning becomes increasingly used in post-training and training costs continue to rise, a data-driven curriculum mechanism that dynamically prioritizes learnable distributions is not just desirable, but necessary.
@@ 2 Background
RL-based LLM Post-training. Reinforcement learning (RL) plays a central role in post-training large language models (LLMs), especially for tasks involving reasoning, subjective preference, or long-horizon control. The RLHF framework  [1, 13, 14, 15, 16] laid the foundation by aligning models using reward signals derived from human preferences. Beyond preference alignment, recent RL-based post-training approaches have notably enhanced LLMs’ capabilities in complex reasoning tasks, particularly coding and mathematics. For instance, RL post-trained model OpenAI o1 [17], o3 [18, 5], DeepSeek-R1 [4] significantly outperform LLMs without RL post-training such as GPT-4o [19] and DeepSeek-V3 [20] on challenging mathematics and coding benchmarks (e.g., AIME [21] and Codeforces [22]). Proximal Policy Optimization (PPO) [23] is widely used in post-training due to its clipped objective, which stabilizes training by preventing large policy updates. PPO remains a strong baseline in many LLM alignment settings. Direct Preference Optimization (DPO) [2] simplifies the pipeline by replacing RL rollouts with a classification-style loss derived from a KL-constrained reward maximization objective. While DPO works well on pairwise preference data, it does not naturally support group-wise or comparative feedback. Group Relative Policy Optimization (GRPO) [3] addresses this limitation by leveraging group-based feedback. For each input prompt $x$, GRPO samples a group of $G$ candidate outputs $\{o_{1},\dots,o_{G}\}\sim\pi_{\text{ref}}(\cdot|x)$ from a frozen reference policy $\pi_{\text{ref}}$. Each output $o_{i}$ is assigned a reward $r_{i}$, and the advantage of $o_{i}$ is computed by normalizing its reward relative to others in the group:
@@
Curriculum Learning for RL. Curriculum learning [24, 25] is a training paradigm that defines a meaningful progression over training samples, typically guiding models to learn from easier examples before progressing to harder ones. In reinforcement learning (RL), task-specific curricula are often used to gradually increase the complexity of the environment [26, 27, 28], aiming to improve the generalization and transferability of learned policies. To reduce reliance on expert-defined task hierarchies, teacher-guided curricula have been proposed [29, 30], where a teacher agent adaptively selects tasks for a student agent, with task selection formalized as a partially observable Markov decision process (POMDP). With the adoption of RL in LLM post-training, curriculum learning has shown potential for improving both training efficiency and model effectiveness. For example, RORL [31] improves alignment by removing the easiest and hardest samples from the training set, keeping only examples of intermediate difficulty. Several recent studies [10, 11, 12, 32] have incorporated curriculum strategies into LLM training. Curri-DPO [10] constructs instance-level curricula by ranking preference pairs based on the score gap between preferred and dispreferred responses, introducing harder pairs gradually during DPO fine-tuning. Kimi k1.5 [11] and Logic-RL [12], on the other hand, use manually defined heuristic curricula with fixed training stages, e.g., models are first trained on “easy” samples for a pre-specified number of steps, then switched to “hard” samples. These strategies rely on static schedules and heuristic difficulty labels, without adapting to the model’s learning progress. R3 [32] addresses sparse reward credit assignment by gradually increasing the generation difficulty—requiring the model to regenerate longer and longer suffixes of demonstration trajectories. While these works demonstrate the benefit of curriculum learning in LLM post-training, most existing approaches focus on instance-level difficulty or use static, manually designed strategies. In contrast, automatic curriculum learning at the distribution level, especially in RL-based post-training, remains underexplored. In this paper, we propose DUMP to fill this gap by adaptively scheduling training over distributions using advantage-based learnability signals.
@@ 3.1 Measuring Learnability via Absolute Advantage
Intuitively, if training on a distribution results in a larger expected advantage magnitude, then that distribution is considered more learnable. The advantage function measures the deviation between an action’s predicted value and its actual return; a large advantage—either positive or negative—indicates that the model’s current policy is still far from optimal on those samples but has a large potential to improve. A small advantage magnitude does not necessarily imply mastery—it may also occur when a task is too difficult or noisy for the model to learn from effectively, resulting in weak or unstable learning signals. To capture this deviation in both directions, we take the absolute value of the advantage. Without this, positive and negative advantages within a batch may cancel out, masking the true extent of the model’s uncertainty or suboptimality. By averaging the absolute advantage over multiple sampled outputs and prompts, we obtain a robust estimate of how much learning signal remains in a given distribution. This expected absolute advantage thus acts as a practical proxy for distribution-level learnability: it reflects how much the model can benefit from training on that distribution. It also has the strength of being lightweight to compute in RL pipelines, as advantage estimates are already generated during rollout.
@@ 4 Experiments and Results
In this section, we first introduce our experiments setup including used models datasets and more implementation details. We then demonstrate the results for the effectiveness of our method DUMP.
@@ 4.1 Experiments Setup
RL Algorithm and LLM Models. We use GRPO [3] as the underlying RL algorithm in our experiments, which is commonly-used in capability-oriented LLM post-training [4]. We use Qwen2.5-7B-Instruct-1M [34] in our experiments. This model is a 7 billion parameter instruction-tuned language model from the Qwen series, which has demonstrated strong long-context capabilities.
Datasets. We evaluate our method on the Knights and Knaves (K&K) puzzle dataset, originally introduced by Xie et al. [35], and widely adopted in RL-based LLM post-training research [12]. This dataset consists of procedurally generated logic puzzles, where each character is either a knight (who always tells the truth) or a knave (who always lies), and the goal is to determine the identity of each character based on their statements. These puzzles require multi-step logical reasoning and are designed to evaluate a model’s capacity for deduction rather than surface-level pattern matching or memorization. An illustration of sample puzzles and the prompting format used during training [12] is provided in Figure 1. The dataset allows precise control over difficulty by varying the number of characters in each puzzle. In our experiments, we synthesize puzzles with character counts ranging from 3 to 14. We treat puzzles with the same number of characters as belonging to the same distribution, resulting in 12 distinct distributions. Each distribution contains 900 training samples and 100 test samples.
Figure 1: Example of puzzle and prompt used in Knights and Knaves (K&K) puzzle dataset.
@@
Figure 2: Effectiveness of DUMP on the K&K puzzle dataset mixed with 12 distributions defined by the number of characters in each puzzle. DUMP consistently achieves higher answer reward on test dataset compared to baseline.
@@
Figure 3: Curriculum (sample counts) induced by DUMP across 12 distributions with increasing difficulty (defined by the number of characters in each puzzle). Simpler distributions are automatically prioritized in early training, while more complex ones are progressively emphasized—both in an entirely automated manner—demonstrating automated distribution scheduling.
  • Format Reward. A binary reward is assigned based on whether the output strictly adheres to the expected format. If the model includes exactly one well-formed <think> and one <answer> section in the correct order, it receives a reward of $+1$; otherwise, it receives a penalty of $-1$.
  • Answer Reward. Conditional on the format being correct, we further evaluate the semantic correctness of the final answer. If the predicted identities fully match the ground truth, the model receives a reward of $+2$; if the answer is incorrect, $-1.5$; and if the answer is missing or unparsable, $-2$.
Other Implementation Details. All experiments are conducted on servers equipped with 8 Nvidia A100 GPUs. Our method is implemented with VeRL [36] LLM Reinforcement Learning framework. We use GRPO [3] as the training algorithm and follow standard practice for actor rollout and optimization. The actor learning rate is set to $1e{-6}$, training batch size is set to 128, and the PPO mini-batch size is 32. KL divergence regularization is applied to encourage alignment with the reference policy, with a KL loss coefficient of 0.001. Each rollout batch contains 16 responses, and we allow for a maximum response length of 20480 tokens during training. The window size $k$ and the temperature $\tau$ in our curriculum learning framework is set to 300 and 0.1, respectively.
@@ 4.2 Effectiveness of DUMP
To evaluate the effectiveness of DUMP in improving post-training efficiency and performance, we compare it against a uniform distribution sampling baseline across 12 distinct data distributions in the K&K puzzle dataset. Each distribution corresponds to a fixed number of characters in the puzzle, ranging from 3 to 14. Figure 2 plots the test answer reward over training steps for each distribution, with and without DUMP. Across all distributions, DUMP consistently outperforms the baseline, achieving faster convergence and higher test performance. The gains are particularly notable in mid- to high-difficulty distributions (e.g., 6 to 12 characters), where uniform sampling tends to struggle due to data underutilization. For example, in the 9-character distribution (Figure 2g), the model trained with DUMP achieves a reward of over 0.5, whereas the baseline remains below 0.0. These results validate the core intuition of DUMP: dynamically adjusting the sampling focus toward high-learnability distributions accelerates policy improvement while avoiding wasted effort on over-saturated or low-signal data. Notably, the improvement is achieved without any curriculum heuristics or manual data ordering—only by observing advantage signals and adapting online.
@@ 4.3 Analyzing the Automated Curriculum by DUMP
To understand how DUMP dynamically allocates training effort across data distributions, we analyze the sampling patterns induced by its UCB-based curriculum mechanism. Figure 3 shows the cumulative number of samples drawn from each distribution (3 to 14 characters) over the course of training. We observe a clear curriculum-like progression: distributions corresponding to simpler puzzles (e.g., 3–5 characters) are heavily sampled in the early stages of training, while more complex distributions (e.g., 10–14 characters) are gradually introduced and increasingly prioritized as training progresses. This pattern aligns with the model’s evolving capacity—early training favors distributions with high initial advantage magnitudes, and as the model saturates on those, DUMP shifts focus to underexplored but learnable distributions. Importantly, this adaptive sampling behavior emerges automatically from empirical advantage signals without requiring manual specification of curriculum order. These results highlight DUMP’s ability to construct an implicit, data-driven curriculum that mirrors traditional easy-to-hard strategies, while remaining responsive to online training dynamics.

### v2

@@ 1 Introduction
Reinforcement learning (RL)-based post-training has emerged as a powerful approach for enhancing the capabilities of large language models (LLMs), particularly in areas requiring structured reasoning, multi-step inference, and task-specific generalization [1, 2, 3, 4]. By leveraging reward signals derived from task performance, human feedback, or domain-specific metrics, RL provides a flexible alternative to supervised fine-tuning. Unlike imitation-based methods that merely mimic reference outputs, RL-based approaches allow models to optimize directly toward behavioral objectives, making them especially effective for boosting model performance on complex reasoning and agentic tasks.
While RL-based post-training has become a key technique for enhancing LLM capabilities in reasoning, alignment, and coding, one foundational challenge remains underexplored: how to dynamically schedule training across heterogeneous data distributions. In practice, LLMs are post-trained on datasets drawn from a wide variety of sources—ranging from factual QA to math problems and coding tasks—each differing in knowledge/capability relevance, and learning difficulty [5, 6, 7]. This heterogeneity is evident in large-scale post-training datasets such as Tülu 3 [7], where prompts span general dialogue, logic puzzles, STEM problems, and multilingual instructions, with widely varying counts, formats, and alignment objectives. More recently, next-generation post-training pipelines (e.g., Seed-Thinking v1.5 [8]) have shifted toward synthetic data generation with controllable parameters—e.g., configuring logical puzzle difficulty. This allows fine-grained control over the data distribution, making distribution-level curriculum learning both feasible and increasingly important. Despite this, most RL-based pipelines still treat all data distributions equally—uniformly sampling tasks throughout training or relying on static, hand-designed curricula. This static treatment ignores the model’s evolving learning needs and underutilizes the training budget. Moreover, it is difficult to handcraft effective curricula when the post-training data comes from multiple distributions lacking clear difficulty labels. As reinforcement learning becomes increasingly used in post-training and training costs continue to rise, a data-driven curriculum mechanism that dynamically prioritizes learnable distributions is not just desirable, but necessary.
@@ 2 Background
RL-based LLM Post-training. Reinforcement learning (RL) plays a central role in post-training large language models (LLMs), especially for tasks involving reasoning, subjective preference, or long-horizon control. The RLHF framework  [1, 12, 13, 14, 15] laid the foundation by aligning models using reward signals derived from human preferences. Beyond preference alignment, recent RL-based post-training approaches have notably enhanced LLMs’ capabilities in complex reasoning tasks, particularly coding and mathematics. For instance, RL post-trained model OpenAI o1 [16], o3 [17, 18], DeepSeek-R1 [4] significantly outperform LLMs without RL post-training such as pre-trained versions of GPT-4o [19] and DeepSeek-V3 [20] on challenging mathematics and coding benchmarks (e.g., AIME [21] and Codeforces [22]). Proximal Policy Optimization (PPO) [23] is widely used in post-training due to its clipped objective, which stabilizes training by preventing large policy updates. PPO remains a strong baseline in many LLM alignment settings. Direct Preference Optimization (DPO) [2] simplifies the pipeline by replacing RL rollouts with a classification-style loss derived from a KL-constrained reward maximization objective. While DPO works well on pairwise preference data, it does not naturally support group-wise or comparative feedback. Group Relative Policy Optimization (GRPO) [3] addresses this limitation by leveraging group-based feedback. For each input prompt $x$, GRPO samples a group of $G$ candidate outputs $\{o_{1},\dots,o_{G}\}\sim\pi_{\text{ref}}(\cdot|x)$ from a frozen reference policy $\pi_{\text{ref}}$. Each output $o_{i}$ is assigned a reward $r_{i}$, and the advantage of $o_{i}$ is computed by normalizing its reward relative to others in the group:
@@
Curriculum Learning for RL. Curriculum learning [24, 25] organizes training by progressing from easy to hard examples. In RL, curricula often follow task complexity [26, 27, 28], or are learned via teacher-student frameworks modeled as partially observable Markov decision process [29, 30]. With the adoption of RL in LLM post-training, curriculum learning has shown potential for improving both training efficiency and model effectiveness. For example, Curri-DPO [9] constructs instance-level curricula by ranking preference pairs based on the score gap between preferred and dispreferred responses, introducing harder pairs gradually during DPO fine-tuning. Kimi k1.5 [10] and Logic-RL [11], on the other hand, use manually defined heuristic curricula with fixed training stages, e.g., models are first trained on “easy” samples for a pre-specified number of steps, then switched to “hard” samples. These strategies rely on static schedules and heuristic difficulty labels, without adapting to the model’s learning progress. While these works demonstrate the benefit of curriculum learning in LLM post-training, most existing approaches focus on instance-level difficulty or use static, manually designed strategies. In contrast, automatic curriculum learning at the distribution level, especially in RL-based post-training, remains underexplored. In this paper, we propose DUMP to fill this gap by adaptively scheduling training over distributions using advantage-based learnability signals.
@@ 3.1 Measuring Learnability via Absolute Advantage
The proof can be found in Appendix A. Intuitively, if training on a distribution results in a larger expected advantage magnitude, then that distribution is considered more learnable. The advantage function measures the deviation between an action’s predicted value and its actual return; a large advantage—either positive or negative—indicates that the model’s current policy is still far from optimal on those samples but has a large potential to improve. A small advantage magnitude does not necessarily imply mastery—it may also occur when a task is too difficult or noisy for the model to learn from effectively, resulting in weak or unstable learning signals. To capture this deviation in both directions, we take the absolute value of the advantage. Without this, positive and negative advantages within a batch may cancel out, masking the true extent of the model’s uncertainty or suboptimality. By averaging the absolute advantage over multiple sampled outputs and prompts, we obtain a robust estimate of how much learning signal remains in a given distribution. This expected absolute advantage thus acts as a practical proxy for distribution-level learnability: it reflects how much the model can benefit from training on that distribution. It also has the strength of being lightweight to compute in RL pipelines, as advantage estimates are already generated during rollout.
@@ 4 Experiments and Results
In this section, we first introduce our experiments setup including used models datasets and more implementation details. We then demonstrate the results for the effectiveness of our method DUMP. More discussion about the comparison to static heuristic curriculum [11, 10] can be found in Appendix C.
@@ 4.1 Experiments Setup
RL Algorithm and LLM Models. We use GRPO [3] as the underlying RL algorithm in our experiments, which is commonly used in capability-oriented LLM post-training [4]. We use Qwen2.5-7B-Instruct-1M [32] and Qwen2.5-3B-Instruct [32] in our experiments.
Datasets and Settings. Multiple datasets are used in our experiments, including Knights and Knaves (K&K) puzzle dataset [33], RuleTaker [34], ProofWriter [35], AR-LSAT [36], LogiQA [37], LogicNLI [38], LongICLBench [39], GSM-8K [40], and AIME 1983-2024 [21]. In our experiments, we consider three different settings. The prompt template used in shown in Figure 3 in the Appendix.
Setting 1: Post-training on K&K puzzles with varying character numbers. The Knights and Knaves (K&K) dataset [33] contains procedurally generated logic puzzles where each character is either a knight (always truthful) or a knave (always lying), and the goal is to infer each character’s identity. The dataset supports fine-grained difficulty control by adjusting the number of characters. We generate puzzles with 3 to 14 characters, treating each character count as a separate distribution—yielding 12 distinct distributions. Each distribution includes 900 training and 100 test samples. We post-train Qwen2.5-7B-Instruct-1M on the combined dataset across all distributions.
Setting 2: Post-training on diverse logic reasoning distributions. We perform post-training using a mixture of logic reasoning datasets, including RuleTaker [34], ProofWriter [35], AR-LSAT [36], LogiQA [37], LogicNLI [38], LongICLBench Geomotion [39], and Knights and Knaves (K&K) [33]. For RuleTaker, ProofWriter, and K&K, we further partition the data distributions by complexity levels: RuleTaker by 2, 3, and 5 required reasoning steps; ProofWriter by 3, 4, and 5 required reasoning steps; and K&K by the number of characters (3–7). In total, we construct 15 logic distributions, each containing 400 training samples. We use Qwen2.5-7B-Instruct-1M for this setting.
Setting 3: Post-training on diverse math reasoning distributions. We also explore post-training on diverse math data. For AIME, we split the data into four distributions based on competition years—1983–1993, 1994–2004, 2005–2015, and 2016–2024—since problem styles evolve significantly over time. We also include GSM-8K as a complementary math dataset. This results in five math distributions in total, with 7473 (GSM-8K), 124, 194, 283, and 238 training samples, respectively. We use Qwen2.5-3B-Instruct for this setting.
  • Format Reward. A binary reward based on whether the output strictly adheres to the expected format. If the model includes exactly one well-formed <think> and one <answer> section in the correct order, it receives a reward of $+1$; otherwise, it receives a penalty of $-1$.
  • Answer Reward. We evaluate the correctness of the final answer. If the predicted identities fully match the ground truth, the model receives a reward of $+2$; if the answer is incorrect, $-1.5$; and if the answer is missing or unparsable, $-2$.
Data Distribution without DUMP with DUMP
RuleTaker 2 Steps 0.79 0.79
RuleTaker 3 Steps 0.76 1.02
RuleTaker 5 Steps 0.56 0.98
ProofWriter 3 Steps 1.18 1.09
ProofWriter 4 Steps 0.97 1.09
ProofWriter 5 Steps 1.24 1.05
AR-LSAT -0.70 -0.52
LogiQA 1.94 1.70
LogicNLI -0.29 -0.23
LongICLBench Geomotion 0.54 0.25
K & K 3 Characters 2.00 2.00
K & K 4 Characters 1.54 1.76
K & K 5 Characters 1.53 1.84
K & K 6 Characters 0.83 1.42
K & K 7 Characters 0.56 1.02
Average 0.90 1.17
Table 1: Test Answer Reward (see Section 4.1) on diverse logic reasoning distributions (Setting 2). The model used here is Qwen2.5-7B-Instruct-1M.
Data Distribution without DUMP with DUMP
GSM-8K 1.50 1.47
AIME 1983-1993 -0.76 -0.39
AIME 1994-2004 -1.50 -1.02
AIME 2005-2015 -0.94 -0.94
AIME 2016-2024 -1.27 -1.27
Average -0.59 -0.43
Table 2: Test Answer Reward (see Section 4.1) on diverse math reasoning distributions (Setting 3). The model used here is Qwen2.5-3B-Instruct.
Other Implementation Details. All experiments are conducted on servers equipped with 8 Nvidia A100 GPUs. Our method is implemented with VeRL [41] LLM Reinforcement Learning framework. We use GRPO [3] as the training algorithm and follow standard practice for actor rollout and optimization. The actor learning rate is set to $1e{-6}$, training batch size is set to 128, and the PPO mini-batch size is 32. KL divergence regularization is applied to encourage alignment with the reference policy, with a KL loss coefficient of 0.001. Each rollout batch contains 16 responses. If not specified, we allow for a maximum response length of 20480 and 4096 tokens during training for Qwen2.5-7B-Instruct-1M and Qwen2.5-3B-Instruct, respectively. The window size $k$ and the temperature $\tau$ in our curriculum learning framework is set to 300 and 0.1, respectively.
@@
Figure 1: Effectiveness of DUMP on the K&K puzzle dataset mixed with 12 distributions defined by the number of characters in each puzzle (Setting 1). DUMP consistently achieves higher answer reward on test dataset compared to baseline. The model used here is Qwen2.5-7B-Instruct-1M.
@@
Figure 2: Curriculum (sample counts) induced by DUMP across 12 K&K puzzle distributions with increasing difficulty defined by the number of characters in each puzzle (Setting 1). Simpler distributions are automatically prioritized in early training, while more complex ones are progressively emphasized—both in an entirely automated manner—demonstrating automated distribution scheduling.
@@ 4.2 Effectiveness of DUMP
Setting 1: Post-training on the combination of K&K puzzle datasets with different number of characters. To evaluate the effectiveness of DUMP in improving post-training efficiency and performance, we compare it against a uniform distribution sampling baseline across 12 distinct data distributions in the K&K puzzle dataset. Each distribution corresponds to a fixed number of characters in the puzzle, ranging from 3 to 14. Figure 1 plots the test answer reward over training steps for each distribution, with and without DUMP. Across all distributions, DUMP consistently outperforms the baseline, achieving faster convergence and higher test performance. The gains are particularly notable in mid- to high-difficulty distributions (e.g., 6 to 12 characters), where uniform sampling tends to struggle due to data underutilization. For example, in the 9-character distribution (Figure 1g), the model trained with DUMP achieves a reward of over 0.5, whereas the baseline remains below 0.0. These results validate the core intuition of DUMP: dynamically adjusting the sampling focus toward high-learnability distributions accelerates policy improvement while avoiding wasted effort on over-saturated or low-signal data. Notably, the improvement is achieved without any curriculum heuristics or manual data ordering—only by observing advantage signals and adapting online.
Setting 2: Post-training on diverse logic reasoning distributions. We apply DUMP to 15 logic reasoning distributions including subsets of RuleTaker, ProofWriter, and K&K (with varying difficulty levels), as well as datasets such as AR-LSAT, LogiQA, LogicNLI, and LongICLBench. As shown in Table 2, DUMP improves the average test answer reward from 0.90 to 1.17. Notable improvements are observed on complex tasks such as AR-LSAT, where the reward increases from -0.70 to -0.52, and K&K 7 Characters, from 0.56 to 1.02. These results demonstrate that DUMP adaptively prioritizes undertrained but learnable distributions, leading to more efficient capability gains.
Setting 3: Post-training on diverse math data distributions. We further evaluate DUMP on GSM-8K and different subsets of AIME grouped by competition years. As shown in Table 2, DUMP raises the average test answer reward from -0.59 to -0.43, with the most significant gain on AIME 1994–2004, where performance improves from -1.50 to -1.02. These results highlight DUMP’s robustness under distribution shifts and data imbalance.
4.3 Ablation Study on the Sampling Strategy
In this section, we ablate the sampling strategy used in DUMP’s UCB-based scheduler. As described in Algorithm 1, our method applies soft sampling controlled by a temperature parameter. The greedy variant (temperature = 0) always selects the distribution with the highest UCB score, while our default uses a small temperature (0.1) to enable probabilistic sampling. We conduct experiments under Setting 1, with a maximum training response length of 10240 tokens. After 100 training steps, the greedy strategy significantly underperforms due to its lack of exploration—it tends to lock onto a single distribution early and fails to adapt. For instance, on the 13- and 14-character K&K tasks, the greedy variant achieves test answer rewards of $-0.91$ and $-1.38$, while soft sampling reaches $-0.66$ and $-1.16$, respectively. These results highlight the importance of maintaining exploration via a non-zero temperature to prevent the scheduler from collapsing onto suboptimal distributions.
@@ 4.4 Analyzing the Automated Curriculum by DUMP
To understand how DUMP dynamically allocates training effort across data distributions, we analyze the sampling patterns induced by its UCB-based curriculum mechanism. Figure 2 shows the cumulative number of samples drawn from each distribution (3 to 14 characters) over the course of training on K&K puzzles with varying character numbers (Setting 1). We observe a clear curriculum-like progression: distributions corresponding to simpler puzzles (e.g., 3–5 characters) are heavily sampled in the early stages of training, while more complex distributions (e.g., 10–14 characters) are gradually introduced and increasingly prioritized as training progresses. This pattern aligns with the model’s evolving capacity—early training favors distributions with high initial advantage magnitudes, and as the model saturates on those, DUMP shifts focus to underexplored but learnable distributions. Importantly, this adaptive sampling behavior emerges automatically from empirical advantage signals without requiring manual specification of curriculum order. These results highlight DUMP’s ability to construct an implicit, data-driven curriculum that mirrors traditional easy-to-hard strategies, while remaining responsive to online training dynamics.